# inkscapelatex=false

### v1

@@ 7.5. GPT-2 Experiments
In this section we report on experiments conducted on the GPT-2 Xl model suggesting that our theoretical findings hold under more general assumptions. Since our results are asymptotic, we need to increase the depth of the GPT-2 XL model. We do so by running the same set of tokens through the model multiple times. In other words, we extract the tokens at the end of the model and feed them to the model for another pass thereby simulating a model of increased length. In total, the tokens pass through the model $2000$ times.
In the first experiment, we prompted the model with:
Describe a futuristic city where humans and robots live together.
Talk about what the city looks like and what daily life is like there.
@@
The experiments were conducted on two different configurations of the GPT-2 XL model, the first using the pre-trained weights provided by the Hugging Face library [34], and the second using randomly generated weight matrices. The multiple passes through the model result in matrices $P$ and $U$ that are time-varying but periodic with period corresponding to the depth of the GPT-2 XL model: 48 layers. To measure how far the tokens are from each other we used the function (25) whose evaluation after each pass is depicted in Figure 6.
Figure 6. Evaluation of the function (25) after each pass through GPT-2 XL with original and random weights.
We can observe that for both configurations the function $E$ decreases with each pass through the model thus implying the tokens converge to a consensus equilibrium. We recall that our theoretical results predict this observation only when feedforward layers are absent. These empirical results suggest that token consensus does not depend on the chosen weight matrices. However, this conclusion is predicated on the matrices $P$ and $U$ being periodic.
@@
To eliminate the periodicity constraint, we repeated the same experiment while randomly generating all the weight matrices with each pass through the model. In addition, we also performed this test by removing the feedforward layers (including the associated normalization function and skip connection) to better understand the impact of these on token consensus. The results are reported in Figure 7 where we can see that convergence towards consensus still occurs.
Figure 7. Comparison between original GPT-2 XL architecture and GPT-2 XL with feedforward layers removed; evaluation of the function (25) after each pass through the model; model weights resampled after each model pass.
In the next experiment we returned to time-varying periodic matrices $P$ and $U$ randomly picked using a Gaussian distribution with zero mean and standard deviation $0.02$. We randomly picked five sentences from [35], presented in Table 2, to be used as prompts and test if consensus depends on the prompt.
Table 2. Five sentences randomly chosen from [35] to be used as prompts.
Prompts
1. A few colleagues and I amused ourselves at a previous IEEE conference on Decision and Control (CDC) with having ChatGPT try to handle all sorts of things, including a failed attempt at having it find Lyapunov function.
2. The instructions that we gave to ChatGPT were, ’’Write a presidential column for the Control Systems Magazine on the future of control,’’ and the resulting column is given verbatim here:
3. Spearheaded by (CSS) Vice-President for Conference Activities Carolyn Beck and Prof. Philip E. Paré, we arrived at a temporary policy, for now, large-scale language models are not allowed as coauthors.
4. Imagine the absurdity of having to agree that ‘‘By uploading this manuscript, I certify that the computations in this submission are made by hand and no calculator was used to perform any of the computations.’’
5. I have no doubt that our warning message will seem as quaint and obsolete as this statement about calculators within a few years.
Since all the prompts have different lengths each prompt was padded with zeros. The results are present in Figure 8 where it can be seen that all tokens tend to converge to a consensus equilibrium for all the tested inputs.
The final experiment was designed to illustrate the deleterious effects of token consensus, using the prompt:
After endless years lost in the shadows of Shakespeare’s sonnets and the melancholic musings of Pessoa, I have glimpsed enlightenment’s elusive light.
Now, on the precipice of my final hour, as the weight of mortality presses upon me, I must reveal to you the one truth that transcends all others—the meaning of life is...
We performed a series of model passes using the pre-trained GPT-2 XL weights and decoded the output, after $1$, $5$, $10$, $100$, and $1000$ passes, using the classifier layer and a greedy sampling approach [^3] By greedy sampling we mean selecting the output with the highest probability.. The results, presented in Table 3, show that as the number of model passes increases, the generated words become increasingly repetitive illustrating the convergence of the tokens to a consensus equilibrium, i.e., model collapse.
Figure 8. Evaluation of the function (25) for the five different prompts in [35].
Table 3. Decoded output of the GPT-2 XL model for different number of passes through the model with original model weights.
# passes Decoded phrase
00 (original prompt) After endless years lost in the shadows of Shakespeare’s sonnets and the melancholic musings of Pessoa, I have glimpsed enlightenment’s elusive light. Now, on the precipice of my final hour, as the weight of mortality presses upon me, I must reveal to you the one truth that transcends all others—the meaning of life is…
$1$ the hours of in the m, the’s playsnets, the worksolic poetryings of therouoa, the finally finallyed the. light light. I I I the eveice of a th exams, I I sun of my begins down me, I am face the the the truth thing that Iends all others:the one of life. lovelove
$5$ ,, of, the last, the, time,, the lastol,ing of the,,, I have beened the. last,. I, I the lastit of the last,, I I lasty the, on me, I have have to you my last and I Ien the the.the one of the. to
$10$ ,, of, the last, the,,,, I last,,ing on the S,, I have beening my. last,, I, I the last, of the,,, I I lasty the, on me, I have have to you my last, I,, the,.I last of the. to to
$100$_ __, in the thermal, the in and_ in in thermal_ in in in thermal __ ___ __ __ ___my in __thermal in_ __the thermal__ the ___in __ __thermal in the in on my, and __ __my the the thermal in in_ in the_ in __thermal_ the_ to __
$1000$___/__/_______________________
__/___/_______/_______________/__\__\___

### v2

@@ 6.2. GPT-2 and GPT-Neo Experiments
In this section we report on experiments conducted on the GPT-2 XL model and the GPT-Neo 2.7B model suggesting that our theoretical findings hold under more general assumptions. Since our results are asymptotic, we need to increase the depth of both models. We do so by running the same set of tokens through the model multiple times. In other words, we extract the tokens at the end of the model, after the final normalization, and feed them to the model for another pass thereby simulating a model of increased length, the code used for our experiments is available at [35].
For all our experiments, we used the same set of 100 random prompts, each generated by uniformly sampling 200 tokens from the GPT-2 tokenizer’s vocabulary. In each experiment, we plot the average of $E$ across all the prompts. As an example, the first 10 tokens of the first sampled prompt are:
divest anxYou coasts Oz
Vi Happy appreciate tcp
In the first experiments, we removed the feedfoward layers of each model, to make them closer to the structure we assume in our theoretical work. The experiments were then repeated without removing the feedforward layers, showing that in both cases convergence to consensus occurs.
Figure 7. Comparison between the GPT-2 XL and GPT-Neo 2.7B architectures with feedforward layers removed; evaluation of the average of Eq.˜25 across all the random prompts.
Figure 8. Comparison between the GPT-2 XL and GPT-Neo 2.7B architectures with the full model; evaluation of the average of Eq.˜25 across all the random prompts.
@@
The experiments were conducted on the standard configuration of the GPT-2 XL model and the GPT-Neo 2.7B model, using the pre-trained weights provided by the Hugging Face library [36]. The multiple passes through the model results in matrices $P$ and $U$ that are time-varying but periodic with period corresponding to the depth each model: 48 layers for the GPT-2 XL and 32 for the GPT-Neo 2.7B. To measure how far the tokens are from each other we used the function Eq.˜25 whose evaluation after each layer is depicted in Figs.˜7 and 8.
We can observe that in both models the average of the function $E$ over all the prompts converge asymptotically to 0, thus implying the tokens converge to a consensus equilibrium. We recall that our theoretical results predict this observation only when feedforward layers are absent, i.e., the case depicted in Fig.˜7. However, as can be seen in Fig.˜8, even when the feedfoward layers are present, convergence still occurs. The rate of convergence appears to be dependent on the weights of the feedfoward layers as their presence increases the convergence rate in the GPT-2 model, but decreases it in the GPT-Neo model. In both cases, these findings suggest that feedfoward layers may not be sufficient to preclude consensus.
@@
Although the previous experimental results suggest that consensus occurs even in the presence of feedfoward layers, it does not address the question of consensus being a structural property of the the transformer architecture or of the choice of weight matrices. To address this question we repeated the experiments by using random matrices in GPT-2 and GPT-Neo. Since this results in time-varying matrices, we further repeated the experiments by randomly selecting new weight matrices before each model pass. Moreover, we conducted these experiments with the full model and also by removing the feedforward layers (including the associated normalization function and skip connection) to better understand the impact of these on token consensus. The results are reported in Figs.˜9 and 10, where we can see that convergence towards consensus still occurs across all experiments. Furthermore, Figs.˜9 and 10 suggest that the feedfoward layer may decrease the convergence rate.
Figure 9. Comparison between the GPT-2 XL and GPT-Neo 2.7B architectures with fixed and randomly chosen weight matrices. Each model was evaluated with and without feedfoward layers using the average of Eq.˜25 across all the random prompts.
Figure 10. Comparison between the GPT-2 XL and GPT-Neo 2.7B architectures with random weight matrices chosen before each model pass. Each model was evaluated with and without feedfoward layers using the average of Eq.˜25 across all the random prompts.
Our experiments suggest that the convergence phenomenon is a product of the structure of the transformers and not of the choice of weights. We observe convergence with trained, random periodic, and random aperiodic matrices $P$ and $U$. In terms of rate of convergence, the choice of weight matrices appears to have an impact with faster convergence being observed when pretrained matrices were used.